OpenStudy (anonymous):

A trapezoid has one base that measures 5 inches and its other base measures 3 inches. What is the area of the trapezoid if its height measures 4 inches? [a] 12 sq. ins. [b] 14 sq. ins. [c] 16 sq. ins. [d] 18 sq. ins.

OpenStudy (danjs):

is it Area = (1/2) * height*(base1 + base2) A = (1/2)*h*(b1 + b2)